OpenStudy (anonymous):

Convert __ (4√x |+3x)/(x^2) to the form ax^n + bx^m

OpenStudy (tkhunny):

1/2 - 2 = -3/2 1 - 2 = -2 You're almost there!

OpenStudy (anonymous):

i need it in the form \[ax^n + bx^m\]

OpenStudy (tkhunny):

Right. I showed you how to find n and m.

OpenStudy (tkhunny):

Whoops! 1 - 2 = -1. Not sure how I managed -2, above.

OpenStudy (anonymous):

would you know what kind of equation/question this is @tkhunny ?

OpenStudy (tkhunny):

?? It's just an algebra exercise. Learn to manipulate and get what you need.

OpenStudy (anonymous):

Thank you. hunny.